On her way back from the amusement park Sally was going 80 miles per hour when she passed a parked highway patrol car. By the time she was half a mile past the patrol car, the officer was chasing her at 100 miles per hour. If these rates remain constant, how long did it take the officer to overtake Sally? Write an equation to represent this situation.
Answer by mananth(16949)

Catch up distance = 0.5 miles
catch up speed = 100-80 mph = 20 mph
Time = d/r
Time = 0.5/20
Time = 1/40 hour
time = 1.5 minutes before the patrol catches up with her